OpenStudy (anonymous):

MEDAL FOR BEST ANSWER The sum of two numbers is Twice the greater number minus the lesser number is 7. What is the greater number? A. 6 B. 7 C. 5 D. 12

OpenStudy (anonymous):

i don't know what the sum is, but you can get your answer with algebra simply make one number 'x' and one number 'y' larger no: x lesser no.: y then form two equations with what you're given x+y=[ungiven sum] --- (1) 2x-y=7 --- (2) for (2), make y the subject as such 2x-y=7 -y=7-2x y=2x-7 then substitute it into (1) to find 'x' x+(2x-7)=[ungiven sum] since x is the larger number, you would've gotten your answer^^

OpenStudy (anonymous):

It help, thank you